The oxidation number of sulfur in H2SO4 is :

A

+2

B

+3

C

+4

D

+6

E

+8

 

Grade:11

1 Answers

Mayank Ranka
3 years ago

Let X be the oxidation number of S in H2SO4
The oxidation  number of hydrogen is +1 and the oxidation number of oxygen is -2.
2(+1)+X+4(2)=02+X8=0X6=0X=+6
Hence, the oxidation number of S in H2SO4is +6.